If n is a positive integer, then the smallest positive integer, n3 + 2n is divisible by

Grade:12th pass

1 Answers

Charchit Tailong
8 years ago
n=0
1
put n=1
1+2=3
n=2
8+4=12
n=3
27+8=35
…...
so the smallest number is 1